I need to find the intersection of two lines
4x+7y= 76
3x-9y= -57
I have tried multiplying them to cancel out X but I cannot find the correct number when solving for y. I got 30.4 but then when I plugged it in I got 72.2 for one and 34.2 for the other, they are supposed to be equal.

Answer by 1335(25)

Instead of multiplying, try using the linear combination method.
Multiply the first equation by 9 and the second equation by 7.
First equation: 36x+63y=684
Second equation: 21x-63y=-399
Now add the two equations together. The Ys will cancel.
57x=285
Now solve for x.
x=5
Plug 5 in for x into either original equation
36(5)+63y=684
Now just solve for y.
180+63y=684
63y=504
y=8
x=5 and y=8